OpenStudy (anonymous):

A swimming pool is being drained. The number of gallons of water in the pool changes with time t according to the equation G = -4t + 100. Which equation correctly solves the given equation for t? A. t = 4G - 100 B.t = -4G + 100 C.t = 100 - G/4 D. t = G - 100/4

OpenStudy (anonymous):

subtract 100 from G first

OpenStudy (anonymous):

G - 100 = -4t

OpenStudy (anonymous):

Then you want to divide by -4

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

-104

OpenStudy (anonymous):

(G-100)/-4 = -4t/-4

OpenStudy (anonymous):

Youre looking for a formula. Not a number

OpenStudy (anonymous):

g=−4t+100

OpenStudy (anonymous):

yes, but they want a formula for t. So we have to solve for t.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

To isolate t, we need to move 100 to the other side which gives us G-100 = -4T. How can we move -4 to the other side of the equation.

OpenStudy (anonymous):

-4*t ?

OpenStudy (anonymous):

We want to undo the multiplication -4*t. What undoes multlipication?

OpenStudy (anonymous):

Division

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Divide both sides by -4. (G-100)/(-4) = -4t/-4t. -4/-4=?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

right so you have (G-100)/(-4) = 1*t

OpenStudy (anonymous):

Do you see the answer?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

We've solved the equation for t

OpenStudy (anonymous):

d

OpenStudy (anonymous):

We can distribute the negative on 4 over G-100 which gives us (100 - G)/4 = t

OpenStudy (anonymous):

b

OpenStudy (anonymous):

no c

OpenStudy (anonymous):

yes c

OpenStudy (anonymous):

thanks